OpenStudy (unicornlovrr):

Which equation is related to 58=x+25 - A. 58+25=x+25-25 B.x=58+25 C.x=33 D.x=83

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)

OpenStudy (festinger):

Move all the numbers to one side of the equation: \[58-25=x\]\[x=33\]

OpenStudy (unicornlovrr):

Thanks can you help me with another one?

OpenStudy (unicornlovrr):

-6 an7/8 ÷ (-3 and 3/4) in simplest form

OpenStudy (festinger):

\[-6\frac{3}{4}\div-3\frac{3}{4}=\frac{27}{4}\times\frac{4}{15}=\frac{9}{5}=1\frac{4}{5}\]